OpenStudy (anonymous):

Given that f(x) = 3x + 1 and g(x) = the quantity of 4x plus 2 divided by 3, solve for g(f(0)).

OpenStudy (anonymous):

2 6 8 9

OpenStudy (anonymous):

The answer would be 2.

OpenStudy (anonymous):

Okay, So. Plug 0 in for X in The F(x) Equation and get 1. Then plug the 1 into the G(x) Equation and get 2. SOOO The overall answer is 2.

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

f(0)=1 so g(1)=(4(1)+2)/3=6/3=2.

OpenStudy (anonymous):

You should really tag me in your questions. xD